OpenStudy (anonymous):

9x^2 +6xy-8y^2 factoring trinomials

OpenStudy (daniel.ohearn1):

Do you know the Fundamental Theorem of Algebra?

OpenStudy (daniel.ohearn1):

You will end up having just two roots. Knowing this you can infer the first term in each root product will be 3x So we have (3x )(3x ) off the bat.

OpenStudy (daniel.ohearn1):

Secondly, what two terms will fit so that if we multiply the each of them with 3x and add each product we 6xy? Both terms must multiply together to get -8y, that should limit it down as well.

Directrix (directrix):

9x^2 +6xy-8y^2 Find two numbers that multiply to 9(-8) and ALSO add to 6 12 and -6 9x^2 +6xy-8y^2 = 9x^2 + 12xy -6xy -8y^2 = 3x(3x + 4y) - 2y(3x + 4y) = Note: (3x + 4y) is a common factor (3x + 4y) (3x - 2y) @Ahinojosa11